A generic solid, X, has a molar mass of 79.9 g/mol. In a constant‑pressure calorimeter, 11.0 g of X is dissolved in 313 g of water at 23.00 °C. X ( s ) ⟶ X ( aq ) The temperature of the resulting solution rises to 26.40 °C. Assume the solution has the same specific heat as water, 4.184 J/(g·°C), and that there is negligible heat loss to the surroundings.

How much heat was absorbed by the solution?

What is the enthalpy of the reaction?
Answer #1

Mass of the solution = 313 + 11.0 = 324 grams

Q = mass * specific heat * ∆T

= 324 * 4.184 * ( 26.40 - 23.00 )

= 4609.1 J

Moles of X = mass / molar mass

= 11.0 / 79.9

= 0.1377

0.1377 moles of X produces = 4609.1 J

1 mole will produce = 4609.1 / 0.1377

= -33472 J/ mol or - 33.472 kj / mol ( heat is released by the reaction)

If you have any questions please comment

If you satisfied with the solution please rate it thanks